# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ 3. Background on Sobolev Spaces
In this section we present some well-known results about Sobolev spaces that will be useful throughout our proofs.
@@ 3.1. Definitions and basic properties
For $1\leq p\leq\infty$, we denote by $\mathrm{W}^{-1,p}(\Omega)$, the dual space of $\mathrm{W}^{1,q}(\Omega)$, where $q$ is the Hölder conjugate of $p$ (i.e., $11/p+1/q=$ 1). More precisely,
@@
Similarly, we denote by $\mathrm{W}^{-1,p}_{0}(\Omega)$ respectively $\mathrm{W}^{-1,p}_{L}(\Omega)$ the dual spaces of $\mathrm{W}^{1,q}_{0}(\Omega)$ respectively $\mathrm{W}^{1,q}_{L}(\Omega)$.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open and connected domain with Lipschitz boundary. Then for any $f\in\mathrm{W}^{1,p}(\Omega)$ with $2\leq p<\infty$ we have
$\lVert f\rVert_{\mathrm{L}^{p}}\lesssim\lVert f\rVert_{\mathrm{H}^{1}}^{1-\theta}\lVert f\rVert_{\mathrm{L}^{2}}^{\theta}\quad\mbox{where}\quad\theta=\frac{d}{2}-\frac{d}{p}.$ (3.2)
@@
For any closed subspace $L\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$, we denote by $\mathrm{H}^{1}_{L}(\Omega)$ the space of Sobolev functions with trace in $L$, i.e.,
$\mathrm{H}^{1}_{L}(\Omega)=\{f\in\mathrm{H}^{1}(\Omega):\gamma f\in L\},$
@@
From the continuity of the trace operator, it is easy to see that $\mathrm{H}^{1}_{L}(\Omega)$ is a closed subspace of $\mathrm{H}^{1}(\Omega)$. Moreover, any right inverse $J:\mathrm{H}^{\frac{1}{2}}(\partial\Omega)\rightarrow\mathrm{H}^{1}(\Omega)$ of the trace operator defines an isomorphism $\mathrm{H}^{1}(\Omega)\cong\mathrm{H}^{1}_{0}(\Omega)\oplus\mathrm{H}^{1/2}(\partial\Omega)$ via the map
@@
From this decomposition, one can show that any closed subspace of $\mathrm{H}^{1}(\Omega)$ containing $C^{\infty}_{c}(\Omega)$ is given by $\mathrm{H}^{1}_{L}(\Omega)$ for a unique $L\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$. More precisely, the following holds.
Proposition 3.9 (Characterization of intermediate spaces of $\mathrm{H}^{1}$).
Let $X\subset\mathrm{H}^{1}(\Omega)$ be a closed subspace such that $C^{\infty}_{c}(\Omega)\subset X$, then there exists a unique closed subspace $L\subset\mathrm{H}^{1/2}(\partial\Omega)$ such that $X=\mathrm{H}^{1}_{L}(\Omega)$.
@@ 3.2. Representation and density on dual Sobolev spaces
$T(f)=\overline{\alpha}\int_{\Omega}f(x)\mathrm{d}x+\int_{\Omega}\overline{V}(x)\bm{\cdot}\nabla f(x)\mathrm{d}x.$ (3.3)
@@
To see that $i$ has closed range, note that if $\{(\int f_{n},\nabla f_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence in ×$\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, then by Poincare’s inequality
@@
Now let $T\in\mathrm{W}^{-1,p}(\Omega)$. As the inverse of $i$ is continuous from $\mathrm{ran}(i)$ to $\mathrm{W}^{1,q}$ by the closed graph theorem, the functional $T\circ i^{-1}$ belongs to the dual of $\mathrm{ran}(i)$. Since $\mathrm{ran}(i)$ is a closed subspace of ×$\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, by the Hahn-Banach theorem there exists a continuous extension of $T\circ i^{-1}$ to ×$\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, denoted here by $T^{\prime}$. Thus, by the Riesz representation theorem in $\mathrm{L}^{q}$ spaces (recall that $1\leq q<\infty$ since $1<p\leq\infty$), there exists $\beta\in\mathbb{C}$ and $V\in\mathrm{L}^{p}(\Omega;\mathbb{C}^{d})$ such that
$T^{\prime}(\beta,W)=\overline{\alpha}\beta+\int_{\Omega}\overline{V}(x)\bm{\cdot}W(x)\mathrm{d}x,\quad\mbox{for any $\beta\in\mathbb{C}$ and $W\in\mathrm{L}^{q}(\Omega;\mathbb{R}^{d})$.}$
@@
$\left|T(f)-\int_{\Omega}\overline{g_{n}(x)}f(x)\mathrm{d}x\right|$ $=\left|\int_{\Omega}\overline{V(x)}\bm{\cdot}\nabla f(x)\mathrm{d}x+\int_{\Omega}\mathrm{div}\,\overline{V_{n}}(x)f(x)\mathrm{d}x\right|=\left|\int_{\Omega}\overline{(V-V_{n})}(x)\bm{\cdot}\nabla f(x)\mathrm{d}x\right|$
@@
Since $\mathrm{W}^{1,q}_{L}(\Omega)$ is a closed subspace of $\mathrm{W}^{1,q}(\Omega)$, any functional in $\mathrm{W}^{-1,p}_{L}(\Omega)$ can be continuously extended to a functional in $\mathrm{W}^{-1,p}(\Omega)$ by the Hahn-Banach theorem. In particular, the representation and density results in Lemmas 3.10 and 3.11 hold for $\mathrm{W}^{-1,p}_{L}(\Omega)$ and $\mathrm{W}^{-1,p}_{0}(\Omega)$ as well.
@@ 3.3. Regularity of reduced densities
We now present a regularity result for reduced densities of wave-functions with finite kinetic energy. This result is a simple application of the GNS, Hölder’s, and Minkowiski integral inequalities.
@@
$\lVert\rho_{\Psi,\Phi}\rVert_{1,p}\lesssim\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}^{1-\frac{d}{p}}\lVert\Phi\rVert_{\mathrm{L}^{2}}^{\frac{d}{p}}+\lVert\Phi\rVert_{\mathrm{H}^{1}}\lVert\Psi\rVert_{\mathrm{H}^{1}}^{1-\frac{d}{p}}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{\frac{d}{p}},\quad\mbox{for any}\quad\begin{cases}1\leq p\leq 2,\quad\mbox{if $d=1$,}\\ 1\leq p<2,\quad\mbox{if $d=2$,}\\ 1\leq p\leq\frac{d}{d-1},\quad\mbox{if $d\geq 3$,}\end{cases}$ (3.5)
@@
for any smooth $\Psi,\Phi$. Then we can apply Minkowiski integral inequality, Hölder’s inequality, and the GNS inequality (3.2) to find that
@@
$\lesssim\int_{\Omega^{\prime}}\lVert\nabla_{x}\Psi\rVert_{\mathrm{L}^{2}_{x}}\lVert\Phi\rVert_{\mathrm{H}^{1}_{x}}^{1-\theta}\lVert\Phi\rVert_{\mathrm{L}^{2}_{x}}^{\theta}+\lVert\nabla_{x}\Phi\rVert_{\mathrm{L}^{2}_{x}}\lVert\Psi\rVert_{\mathrm{H}^{1}_{x}}^{1-\theta}\lVert\Psi\rVert_{\mathrm{L}^{2}_{x}}^{\theta}\mathrm{d}y$
$\lesssim\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}^{1-\theta}\lVert\Phi\rVert_{\mathrm{L}^{2}}^{\theta}+\lVert\Phi\rVert_{\mathrm{H}^{1}}\lVert\Psi\rVert_{\mathrm{H}^{1}}^{1-\theta}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{\theta},$
@@
where $\theta=d(\frac{1}{2}-\frac{p-2}{2p})=\frac{d}{p}$. As this holds for any smooth $\Psi,\Phi$, the result follows by an approximation argument. The restriction $p<2$ for $d=2$ is necessary in (3.5) because the GNS inequality (3.2) is not valid in the end point case $(p,d)=(\infty,2)$. On the other hand, the case $(p,d)=(\infty,1)$ is allowed in (3.2) and therefore $p=2$ in (3.5) is allowed for $d=1$. ∎
@@
An important feature of inequality (3.5) is that only the dimension of the set $\Omega$ is relevant for the regularity of the reduced densities. Moreover, in the case where $d^{\prime}=0$, i.e., we consider only $\Psi,\Phi\in\mathrm{H}^{1}(\Omega)$ and $\rho_{\Psi,\Phi}=\overline{\Psi}\Phi$, estimate (3.5) reduces to the standard (Sobolev) regularity expected by the product of two $\mathrm{H}^{1}$ functions in $\mathbb{R}^{d}$. In particular, it implies that the set $\mathrm{H}^{1}(\Omega)$ for an interval $\Omega=(a,b)\subset\mathbb{R}$ is an algebra of functions. This property was crucial in [Cor25, SPR+24] but will play no special role here.
4. Schrödinger operators with distributional potentials
In this section, we present the construction of self-adjoint realizations of $H_{N}(v,w)$ in detail. We also present some important examples of distributional potentials for which this construction holds. For convenience of the reader, we recall the basic results on the theory of quadratic forms and self-adjoint extensions in Appendix A.
4.1. Self-adjoint realizations
$|v(\rho_{\Psi,\Psi})+w(\rho_{2,\Psi,\Psi})|\leq\epsilon\lVert\Psi\rVert_{\mathrm{H}^{1}}^{2}+\frac{C}{\epsilon}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{2},\quad\mbox{for any $\Psi\in\mathrm{H}^{1}(I_{N})$,}$ (4.1)
and any $\epsilon>0$ for some constant $C=C(v,w)>0$. Hence, from the celebrated KLMN theorem (cf. Theorem A.5), for any closed subspace $L\subset\mathrm{H}^{1/2}(\partial\Omega)$, we can construct a unique self-adjoint operator associated to the sesquilinear form
$a_{v,w}(\Psi,\Phi)=\int_{I_{N}}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x+v(\rho_{\Psi,\Phi})+w(\rho_{2,\Psi,\Phi}),\quad\mbox{for $\Psi,\Phi\in\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$.}$ (4.2)
Precisely, the following holds.
Lemma 4.1 (Schrödinger operators with distributional potentials).
Let $N\in\mathbb{N}$, $v\in\mathcal{V}$, and $w\in\mathcal{W}$. Moreover, let $L\subset\mathrm{H}^{1/2}(\partial I_{N})$ be a closed subspace and $\mathrm{H}^{1}_{L}(I_{N})$ the trace-restricted Sobolev space introduced in Definition 3.8. Then the quadratic form in (4.2) is closed, symmetric and semibounded. In particular, there exists a unique self-adjoint operator $H_{N}^{L}(v,w)$ associated to this form. Moreover, $H_{N}^{L}(v,w)$ is semibounded and has purely discrete spectrum.
Remark 4.2 (Notation).
For the rest of the paper, we shall denote by $H_{N}(v,w;\Gamma)$ the self-adjoint realization of $a_{v,w}$ with form domain $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ and by $H^{\alpha}_{N}(v;w;\Gamma)$ the realization with form domain $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$.
4.2. From single-particle to many-particle boundary conditions
To this end, we start with the following lemma.
Lemma 4.3 (Many-particle form domain).
Let $\alpha\in\mathbb{R}$ and $\mathrm{H}^{1}_{\alpha}(I)$ be the space introduced in (3.5). Then for any $N\in\mathbb{N}$, the span of Slater determinants
is dense in $\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$.
Proof.
We first observe that the projection on the antisymmetric space, explictly given by
$\Phi\mapsto(P_{N}\Phi)(x)=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\operatorname{sign}(\sigma)\Phi(\sigma x)$
is continuous with respect to the $\mathrm{H}^{1}$ topology and maps $N$-fold tensor products to $N$-fold Slater determinants. Therefore, it suffices to show that the space
$\alpha\partial_{x}\varphi_{j}(0)-\partial_{x}\varphi_{j}(1)=0.$
$=\sum_{k=1}^{N}\lambda_{j_{k}}\langle\Psi,\varphi_{j_{1}}\otimes...\otimes\varphi_{j_{k}}\rangle_{\mathrm{L}^{2}}+\langle\Psi,\varphi_{j_{1}}\otimes...\otimes\varphi_{j_{N}}\rangle_{\mathrm{L}^{2}}$
As a consequence of the previous lemma, we have the following relation between the single-particle operator $h_{\alpha}(v)=-\Delta_{\alpha}+v$ and the $N$-particle operator $H_{N}^{\alpha}(v,0)$ associated to the form
$a_{v}(\Psi,\Psi)=\int_{I_{N}}|\nabla\Psi(x)|^{2}\mathrm{d}x+v(\rho_{\Psi}),\quad\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}.$
Lemma 4.4 (From single-particle to non-interacting many-particle systems).
Let $v\in\mathcal{V}$, $\alpha\in\mathbb{C}^{2}$, and $N\in\mathbb{N}$. Let $\{\psi_{j}\}_{j\in\mathbb{N}}$ be an orthogonal basis of eigenfunctions of $h_{\alpha}(v)=-\Delta_{\alpha}+v$ with corresponding eigenvalues $\{\lambda_{j}\}_{j\in\mathbb{N}}$ ordered in non-decreasing order. Then the Slater determinants
$\{\Psi_{J}=\psi_{j_{1}}\wedge...\wedge\psi_{j_{N}}:J=\{j_{1},...,j_{N}\}\in\mathbb{N}^{N}\quad\mbox{such that}\quad 1\leq j_{1}<j_{2}...<j_{N}\}$ (4.3)
Proof.
From a long but straightforward calculation we find that
$a_{v}(\Psi_{J},\varphi_{1}\wedge...\wedge\varphi_{N})=\lambda_{J}\langle\Psi_{J},\varphi_{1}\wedge...\wedge\varphi_{N}\rangle_{\mathrm{L}^{2}},\quad\mbox{for any $\{\varphi_{j}\}_{j=1}^{N}\subset\mathrm{H}^{1}_{\alpha}(I)$,}$
where $\Psi_{J}=\psi_{j_{1}}\wedge...\wedge\psi_{j_{N}}$ and $\lambda_{J}=\sum_{k=1}^{N}\lambda_{j_{k}}$. Thus by Lemma 4.3 and approximation we have
$a_{v}(\Psi_{J},\Phi)=\lambda_{J}\langle\Psi_{J},\Phi\rangle\quad\mbox{for any $\Phi\in\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$.}$
4.3. Examples of distributional potentials
To illustrate how large the class of admissible external and interaction potentials in Lemma 4.1 is, let us present a few important examples.
  1. (1)
  2. (2)
    The $\delta$-interaction potential $w=\delta(x-y)$, defined via
    $w_{\delta}(\rho_{2,\Psi,\Phi})=\int_{I}\rho_{2,\Psi,\Phi}(x,x)\mathrm{d}x.$
  3. (3)
    belongs to $\mathrm{W}^{-1,q}(I_{2})$.
  4. (4)
    Lemma 4.1 can also be extended to the case of $k^{th}$-body distributional potentials for $k\geq 3$. To be precise, for any $w_{k}\in\mathrm{W}^{-1,q}(I_{k})$ with $q>k$, one can use Lemma 3.13 and Young’s inequality to show that the form $(\Psi,\Phi)\mapsto w_{k}(\rho_{k,\Psi,\Phi})$, where
    $\rho_{k,\Psi,\Phi}(x_{1},..,x_{k})=\frac{N!}{(N-k)!k!}\int_{I_{N-k}}\overline{\Psi(x_{1},...,x_{N})}\Phi(x_{1},...,x_{N})\mathrm{d}x_{k+1}...\mathrm{d}x_{N},$
    is $\Delta$-bounded with relative bound 00. Therefore, any Schrödinger operator of the form
    is also included for $N\geq 6$.

### v2

@@ 3. Background on Sobolev Spaces
In this section we recall some well-known results about Sobolev spaces that will be used throughout our proofs.
@@ 3.1. Definitions and basic properties
For $1\leq p\leq\infty$, we denote by $\mathrm{W}^{-1,p}(\Omega)$, the dual space of $\mathrm{W}^{1,q}(\Omega)$, where $q$ is the Hölder conjugate of $p$ (i.e., $1/p+1/q=1$). More precisely,
@@
Similarly, we denote by $\mathrm{W}^{-1,p}_{0}(\Omega)$ the dual spaces of $\mathrm{W}^{1,q}_{0}(\Omega)$.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open and connected domain with Lipschitz boundary. Then for any $f\in\mathrm{W}^{1,p}(\Omega)$ with $2\leq p<\infty$ such that $\theta=d/2-d/p\in[0,1]$ we have
$\lVert f\rVert_{\mathrm{L}^{p}}\lesssim\lVert f\rVert_{\mathrm{H}^{1}}^{\theta}\lVert f\rVert_{\mathrm{L}^{2}}^{1-\theta}\quad\mbox{where}\quad\theta=\frac{d}{2}-\frac{d}{p}.$ (3.2)
@@
For any closed subspace $B\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$, we denote by $\mathrm{H}^{1}_{B}(\Omega)$ the space of Sobolev functions with trace in $B$, i.e.,
$\mathrm{H}^{1}_{B}(\Omega)=\{f\in\mathrm{H}^{1}(\Omega):\gamma f\in B\},$
@@
From the continuity of the trace operator, it is easy to see that $\mathrm{H}^{1}_{B}(\Omega)$ is a closed subspace of $\mathrm{H}^{1}(\Omega)$. Moreover, any right inverse $J:\mathrm{H}^{\frac{1}{2}}(\partial\Omega)\rightarrow\mathrm{H}^{1}(\Omega)$ of the trace operator defines an isomorphism $\mathrm{H}^{1}(\Omega)\cong\mathrm{H}^{1}_{0}(\Omega)\oplus\mathrm{H}^{1/2}(\partial\Omega)$ via the map
@@
From this decomposition, it is easy to see that any closed subspace of $\mathrm{H}^{1}(\Omega)$ containing $C^{\infty}_{c}(\Omega)$ is given by $\mathrm{H}^{1}_{B}(\Omega)$ for a unique closed subspace $B\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$. More precisely, the following holds.
Proposition 3.9 (Intermediate spaces between $\mathrm{H}^{1}_{0}$ and $\mathrm{H}^{1}$).
Let $X\subset\mathrm{H}^{1}(\Omega)$ be a closed subspace such that $C^{\infty}_{c}(\Omega)\subset X$, then there exists a unique closed subspace $B\subset\mathrm{H}^{1/2}(\partial\Omega)$ such that $X=\mathrm{H}^{1}_{B}(\Omega)$.
@@ 3.2. Representation and density on dual Sobolev spaces
$T(f)=\overline{\alpha}\int_{\Omega}f(x)\mathrm{d}x+\int_{\Omega}\overline{V}(x)\boldsymbol{\cdot}\nabla f(x)\mathrm{d}x.$ (3.3)
@@
To see that $i$ has closed range, note that if $\{(\int f_{n},\nabla f_{n})\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, then by Poincare’s inequality
@@
Now let $T\in\mathrm{W}^{-1,p}(\Omega)$. As the inverse of $i$ is continuous from $\mathrm{ran}(i)$ to $\mathrm{W}^{1,q}$ by the closed graph theorem, the functional $T\circ i^{-1}$ belongs to the dual of $\mathrm{ran}(i)$. Since $\mathrm{ran}(i)$ is a closed subspace of $\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, by the Hahn-Banach theorem there exists a continuous extension of $T\circ i^{-1}$ to $\mathbb{C}\times\mathrm{L}^{q}(\Omega;\mathbb{C}^{d})$, denoted here by $T^{\prime}$. Thus, by the Riesz representation theorem in $\mathrm{L}^{q}$ spaces (recall that $1\leq q<\infty$ since $1<p\leq\infty$), there exists $\beta\in\mathbb{C}$ and $V\in\mathrm{L}^{p}(\Omega;\mathbb{C}^{d})$ such that
$T^{\prime}(\beta,W)=\overline{\alpha}\beta+\int_{\Omega}\overline{V}(x)\boldsymbol{\cdot}W(x)\mathrm{d}x,\quad\mbox{for any $\beta\in\mathbb{C}$ and $W\in\mathrm{L}^{q}(\Omega;\mathbb{R}^{d})$.}$
@@
$\left|T(f)-\int_{\Omega}\overline{g_{n}(x)}f(x)\mathrm{d}x\right|$ $=\left|\int_{\Omega}\overline{V(x)}\boldsymbol{\cdot}\nabla f(x)\mathrm{d}x+\int_{\Omega}\mathrm{div}\,\overline{V_{n}}(x)f(x)\mathrm{d}x\right|=\left|\int_{\Omega}\overline{(V-V_{n})}(x)\boldsymbol{\cdot}\nabla f(x)\mathrm{d}x\right|$
@@
Since $\mathrm{W}^{1,q}_{B}(\Omega)$ is a closed subspace of $\mathrm{W}^{1,q}(\Omega)$, any functional in $\mathrm{W}^{-1,p}_{B}(\Omega)$ can be continuously extended to a functional in $\mathrm{W}^{-1,p}(\Omega)$ by the Hahn-Banach theorem. In particular, the representation and density results in Lemmas 3.10 and 3.11 hold for $\mathrm{W}^{-1,p}_{B}(\Omega)$ and $\mathrm{W}^{-1,p}_{0}(\Omega)$ as well.
@@ 3.3. Regularity of reduced densities
We now present a regularity result for reduced densities of wave-functions with finite kinetic energy. This result is a simple application of the GNS, Hölder’s, and Minkowski integral inequalities.
@@
$\lVert\rho_{\Psi,\Phi}\rVert_{1,p}\lesssim\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}^{d-\frac{d}{p}}\lVert\Phi\rVert_{\mathrm{L}^{2}}^{1-d+\frac{d}{p}}+\lVert\Phi\rVert_{\mathrm{H}^{1}}\lVert\Psi\rVert_{\mathrm{H}^{1}}^{d-\frac{d}{p}}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{1-d+\frac{d}{p}},\quad\mbox{for any}\quad\begin{cases}1\leq p\leq 2,\quad\mbox{if $d=1$,}\\ 1\leq p<2,\quad\mbox{if $d=2$,}\\ 1\leq p\leq\frac{d}{d-1},\quad\mbox{if $d\geq 3$,}\end{cases}$ (3.5)
@@
for any smooth $\Psi,\Phi$. Then we can apply Minkowski integral inequality, Hölder’s inequality, and the GNS inequality (3.2) to find that
@@
$\lesssim\int_{\Omega^{\prime}}\lVert\nabla_{x}\Psi\rVert_{\mathrm{L}^{2}_{x}}\lVert\Phi\rVert_{\mathrm{H}^{1}_{x}}^{\theta}\lVert\Phi\rVert_{\mathrm{L}^{2}_{x}}^{1-\theta}+\lVert\nabla_{x}\Phi\rVert_{\mathrm{L}^{2}_{x}}\lVert\Psi\rVert_{\mathrm{H}^{1}_{x}}^{\theta}\lVert\Psi\rVert_{\mathrm{L}^{2}_{x}}^{1-\theta}\mathrm{d}y$
$\lesssim\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}^{\theta}\lVert\Phi\rVert_{\mathrm{L}^{2}}^{1-\theta}+\lVert\Phi\rVert_{\mathrm{H}^{1}}\lVert\Psi\rVert_{\mathrm{H}^{1}}^{\theta}\lVert\Psi\rVert_{\mathrm{L}^{2}}^{1-\theta},$
@@
where $\theta=d(\frac{1}{2}-\frac{2-p}{2p})=d(1-\frac{1}{p})$. As this holds for any smooth $\Psi,\Phi$, the result follows by an approximation argument. The restriction $p<2$ for $d=2$ is necessary in (3.5) because the GNS inequality (3.2) is not valid in the end point case $(p,d)=(\infty,2)$. On the other hand, the case $(p,d)=(\infty,1)$ is allowed in (3.2) and therefore $p=2$ in (3.5) is allowed for $d=1$. ∎
@@
An important feature of inequality (3.5) is that only the dimension of the set $\Omega$ is relevant for the regularity of the reduced densities. Moreover, in the case where $d^{\prime}=0$, i.e., we consider only $\Psi,\Phi\in\mathrm{H}^{1}(\Omega)$ and $\rho_{\Psi,\Phi}=\overline{\Psi}\Phi$, estimate (3.5) reduces to the standard (Sobolev) regularity expected by the product of two $\mathrm{H}^{1}$ functions in $\mathbb{R}^{d}$. In particular, it implies that the set $\mathrm{H}^{1}(\Omega)$ for an interval $\Omega=(a,b)\subset\mathbb{R}$ is an algebra of functions. This property was crucial in [Cor25a, SPR+24] but will play no special role here.